x divided by 5 minus 5 equals 1
x/5 - 5 = 1

Answer by ehollins1(8)

x/5-5=1 cancel out the -5 to get x/5 alone and add the 5 to the one like so
x/5-5+5=1+5 now you are left with the following:
x/5=6 now simple fractions come into play to get the x alone by multiplying 5 on both sides like so
5/1*x/5=6*5 now u are left with the following:
x=6*5
solution x=30
****check solution****
x=30
x/5-5=1
30/5-5=
6-5=1